I always have problems with this type of word problem How can I solve this and write the equation for it? please help
The length of a rectangular playing field is 5 ft less than twice its width. If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by ankor@dixie-net.com(22740)

he length of a rectangular playing field is 5 ft less than twice its width. If the perimeter of the playing field is 230 ft, find the length and width of the field.
:
Let x = width
:
It says
"the length of a rectangular playing field is 5 ft less than twice its width."
Therefore:
length = (2x - 5)
;
If the perimeter of the playing field is 230 ft,
Perimeter: 2L + 2W = 230
:
Substitute (2x-5) for L and x for w and you have:
2(2x-5) + 2x = 230
:
WE can simplify, divide each term by 2 and you have
2x - 5 + x = 115
2x + x = 115 + 5; add 5 to both sides:
3x = 120
x = 120/3
x = 40 ft is the width:
:
Find the length; L = 2W-5
2(40) - 5 = 75 ft is the length
:
:
Check solution; find the perimeter with these values:
2(75) + 2(40) = 230; confirms our solution
:
;
Make sense to you??